Mp025 Aom Settings

Light from MP023

Light is picked off from the K repump line in MP023, after it has been locked to the 39K F=2 -> 2P3/2 (unresolved), then shifted an additional ~-308MHz by AOM1 and AOM2:

AOM1 1x(-116MHz)
AOM2 2x(-96MHz)
Total -308MHz

note: values on MP023 Laser Settings may be a little different.


K Level Diagram

Kleveldiagram

**40K Trapping
MOT / Imaging (addressing F = 2 -> F' = 3)

AOM Name AOM Frequency (MHz) Order Number Passes Actual Optical Freq Shift (MHz)
AOM1 (sat. spec.) 65 +1 2 65
AOM2 (freq. control) 86 (MOT) / 70 (imaging) -1 2 -172 / -140
AOM3 (amp. control) 65 +1 1 80
Locking Transition F = 2 -> F' = 3 Total Shift -27 / +5
Unless otherwise stated, the content of this page is licensed under Creative Commons Attribution-ShareAlike 3.0 License